OpenStudy (anonymous):

What is the value of the x-coordinate of the vertex of the function shown below? f(x)=(x+4)(x-2)

OpenStudy (anonymous):

The vertex is halfway between the two 0's

OpenStudy (anonymous):

Can you easily identify the 2 values for x that will make f(x) = 0

OpenStudy (anonymous):

If not you can alternatively re-write it in vertex form

OpenStudy (anonymous):

Thoughts? Questions?

OpenStudy (anonymous):

so i just plug in zero for x and zer ofor y and then i find the two zeros and the vertex is between them

OpenStudy (anonymous):

f(x)=(x+4)(x-2) -> If x is -4 or 2 in the above, then f(x) = 0

OpenStudy (anonymous):

No, just plug in 0 for y. Find the 2 x values.

OpenStudy (anonymous):

Then find the midpoint between those 2 x values.

OpenStudy (anonymous):

The only to values of x for which the expression (x+1)(x-2) is zero, is either when x=-1 or when x=2

OpenStudy (anonymous):

correction - typo - I meant "4" above , not "1" .

OpenStudy (anonymous):

so now, all that is left is to find the midpoint between -4 and 2

OpenStudy (anonymous):

I think we need a graph http://www.wolframalpha.com/input/?i=f%28x%29%3D%28x%2B4%29%28x-2%29 U can see where the zeros are at -4 and 2. the x coordinate is midway between them. (In case it is confusing u, you can just think of "f(x)" as meaning "y")

OpenStudy (anonymous):

so the answer is 1,-5 like thats the vertex and then the real answer is just 1

OpenStudy (anonymous):

I think you're being confused a bit by all the different info. In general, do you know of any ways to find the vertex?

OpenStudy (anonymous):

its p+q/2 which would be 4+-2/2 so 1

OpenStudy (anonymous):

It's -1 actually. -4 is farther negative than 2 is positive, so the mid point between them must be negative also. The distance between them is |-4 - 2| = |-6| = 6 and half that distance is 3, and 3 more than -4 is -1. So the vertex is at -1.

OpenStudy (anonymous):

Question What is the value of the x-coordinate of the vertex of the function shown below? f(x)=(x+4)(x-2) Responses The vertex is halfway between the two 0's (Polpak) f(x)=(x+4)(x-2) -> If x is -4 or 2 in the above, then f(x) = 0 (me) I think we need a graph http://www.wolframalpha.com/input/?i=f%28x%29%3D%28x%2B4%29%28x-2%29 (me) It's -1 actually. -4 is farther negative than 2 is positive, so the mid point between them must be negative also. The distance between them is |-4 - 2| = |-6| = 6 and half that distance is 3, and 3 more than -4 is -1. So the vertex is at -1.(Polpak) All that's left to be said is plug the -1 into f(x) to get the y value (-1+4)(-1-2)=-9